OpenStudy (anonymous):

An oil-drilling company knows that it costs $25,000 to sink a test well. If oil is hit, the income for the drilling company will be $445,000. If only natural gas is hit, the income will be $155,000. If nothing is hit, there will be no income. If the probability of hitting oil is 1/40 and if the probability of hitting gas is 1/20, what is the expectation for the drilling company?

OpenStudy (tkhunny):

P(Test Well) = 1.00 Thus: 1.00*(-25,000) P(Gas) = 1/20 Thus: (1/20)*(155,000) P(Oil) = 1/40 Thus: (1/40)*(445,000) Those are the three pieces. Add them up.